OpenStudy (anonymous):

(3.1 × 10^11) ÷ (6.2 × 10^6) can anyone explain

OpenStudy (anonymous):

Solve by first doing your exponents, then whats inside the prentisis, then divide the 2

OpenStudy (gw2011):

The answer is 5x10^4. This is how you get the answer: 10^11-^6=10^5 (since the base of 10 is the same in the numerator and denominator, then you subtract the denominator exponent 6 from the numerator exponent 11). Now you divide: 3.1/6.2=0.5 Combining the above two operations you get: 0.5x10^5. Now move the decimal point of 0.5 one place to the right to get: 5x10^4 (this is the accepted way of writing this type of number, i.e., 0.5 is written as 5 or 5.0). You can verify this answer by converting (3.1x10^11) to: 310,000,000,000 (you had to move the decimal point 11 places to the right) Then convert (6.2x10^6) to get: 6,200,000 When you divide these two long numbers you will get 5x10^4. The only reason I converted these numbers to the long form is to prove the correctnes of the answer of 5x10^4. You can see that using 10^to a power is more convenient and easier to solve than the long way. Using powers of 10 is the scientific notation.